### v1

@@ 4 Sweeping
    Take a loop: Let $\alpha$ be a curve that does not intersect $\gamma$ and such that $\alpha$ is visible from $\gamma$ via the visibility arc $\tau(u,v)$. We define “taking a loop” as the operation that modifies $\gamma$ to $\gamma^{\prime}$ as follows: Taking two points $u^{\prime},u^{\prime\prime}$ very close to $u$ on either side to $u$ on $\gamma$, we replace the segment of $\gamma$ between $u^{\prime}$ and $u^{\prime\prime}$ by a curve between $u^{\prime}$ and $u^{\prime\prime}$ that lies very close to $\tau(u,v)$ and loops around $v$ crossing $\alpha$ twice. See Figures 1(a) and 1(b). More formally, $\gamma$ is modified to $\gamma^{\prime}$ so that $\tilde{\gamma^{\prime}}=\tilde{\gamma}\setminus(\tau(u,v)\oplus B_{\epsilon})$ where $\oplus$ denotes Minkowski sum and $B_{\epsilon}$ is a ball of radius $\epsilon$ for an arbitrarily small $\epsilon$. In the rest of the paper, we avoid such formal definitions and resort to figures for ease of exposition. However, our informal definitions using figures can easily be formalized.
@@
    Bypass a digon cell: Suppose that a curve $\alpha$ and the sweep curve $\gamma$ form digon cell $D$ with vertices $u$ and $v$. Then, we define the operation of “bypassing $D$” as the modification of $\gamma$ to a curve $\gamma^{\prime}$ that goes around the digon cell $D$ as shown in Figures 2(a) and 2(b).
@@
    Bypass a triangle cell: Let $\alpha$ and $\beta$ be two curves that along with the sweep curve $\gamma$ define a triangle cell $T$. We define the operation of “bypassing $T$” as the modification of $\gamma$ to a curve $\gamma^{\prime}$ which goes around $T$ as show in Figures 3(a) and 3(b).
@@
    Bypass a visible vertex: Let $v$ be a vertex lying inside $\tilde{\gamma}$ where the curves $\alpha$ and $\beta$ intersect. If there is a visibility arc $\tau(u,v)$ from a point $u$ on $\gamma$ to $v$ whose interior lies in the interior of $\tilde{\alpha}$ as well as $\tilde{\beta}$, then by definition, the interior of $\tau(u,v)$ also lies in the interior of $\tilde{\gamma}$. In this case, we refer to $\tau(u,v)$ as the bypassability curve for $v$. “Bypassing $v$” is the modification of $\gamma$ to a curve $\gamma^{\prime}$ that takes a loop around $v$ as shown in Figures 4(a) and 4(b). Note that unlike the previous operations, the orientation of the curves are important here.
@@ 7 A simpler proof of the theorem of Snoeyink and Hershberger
Let $\Gamma$ be an arrangement of 2-intersecting curves with sweep curve $\gamma$, and let $P$ be a specified point in $\tilde{\gamma}$. Our goal is to sweep the arrangement of $\Gamma$ with $\gamma$ so that the arrangement remains 2-intersecting throughout the process and $P$ remains inside $\tilde{\gamma}$ until no curve in $\Gamma\setminus\gamma$ intersects $\tilde{\gamma}$. We do not use the operation of bypassing a visible vertex, as this violates the 2-intersecting property of the curves. We restrict ourselves to the remaining sweeping operations, namely: $(i)$ taking a loop on a curve that lies entirely in $\tilde{\gamma}$, $(ii)$ bypassing a digon cell, and $(iii)$ bypassing a triangle cell.
@@
Broad idea. In the restricted setting where each pair of curves in $\Gamma\setminus\{\gamma\}$ intersects at most once inside $\tilde{\gamma}$, we can show that the arrangement is sweepable. Therefore, a simplest non-sweepable arrangement must have at least one pair of curves that intersect twice (i.e., they form a digon) in $\tilde{\gamma}$. In this case, we carefully modify the arrangement by bypassing a minimal digon or a minimal triangle (defined in Section 5) to obtain a simpler non-sweepable arrangement, thus arriving at a contradiction.
@@
Let $\Gamma$ be a set of 2-intersecting curves with sweep curve $\gamma$ s.t. all curves in $\Gamma\setminus\{\gamma\}$ intersect $\gamma$, and any pair of curves in $\Gamma\setminus\{\gamma\}$ intersect at most once in $\tilde{\gamma}$. Then, there is a sweepable cell on $\gamma$, i.e., there is either a digon/triangle cell on $\gamma$ that does not contain $P$.
In order to prove the above lemma, we need the following claim which though not explicitly stated is also proved as part of Lemma 3.1 of [23].
@@
Proof 7.1.
We prove by induction on the number of curves intersecting the side of $T$ on $\gamma$. If no curve intersects the interior of this side, since the curves are 1-intersecting in $\tilde{\gamma}$, it follows that $T$ is a triangle-cell. Assume that the claim holds for any half-triangle with less than $k$ curves (where $k\geq 1$) intersecting the side of $T$ on $\gamma$. Suppose now that there are $k$ curves intersecting the interior of the side of $T$ on $\gamma$.
Let $x$ denote the intersection point of $\beta$ and $\gamma$ on the boundary of $T$. Let $v$ denote the intersection point of $\alpha$ and $\beta$ on the boundary of $T$. For every curve $\delta^{\prime}$ intersecting the interior of the side $T$ on $\gamma$, $\delta^{\prime}$ intersects it at one point, and $\delta^{\prime}$ and $\gamma$ intersect at two points. Therefore, since $T$ is a half-triangle with edge $\alpha$, $\delta^{\prime}$ intersects the side $T$ on $\beta$. Walking from $x$ to $v$ along $\beta$, let $\delta$ be the first curve intersecting $\beta$. The curve $\delta$ intersects the interior of the side of $T$ on $\gamma$. Now, $\beta$ and $\delta$ form a half-triangle $T^{\prime\prime}$ on $\gamma$ with edge $\beta$ and s.t. less than $k$ curves intersect the interior of the side of $T^{\prime\prime}$ on $\gamma$. Since every curve in $\Gamma$ has at most one intersection point on the side of $T$ on $\gamma$, the same holds for $T^{\prime\prime}$. Hence, by the inductive hypothesis, there is a triangle cell $T^{\prime}$ on $\gamma$ that lies in $T^{\prime\prime}$. Since $T^{\prime\prime}$ lies in $T$, $T^{\prime}$ is the claimed triangle cell in $T$.
Proof 7.2 (Proof of Lemma 8).
@@
The proof of the main result in this section (Theorem 7.3) follows the structure of Theorem 1, i.e., proof for non-piercing regions. Unlike in the non-piercing case however, since the curves are 2-intersecting, no pair of curves form a negative lens. Let $L$ be a minimal lens in a simplest non-sweepable arrangement $\Gamma$ defined by curves $\alpha,\beta\in\Gamma$. We split the proof into two cases: In the first case, there is a curve $\delta\in\Gamma\setminus\{\alpha,\beta\}$ intersecting $L$. In this case, we show that we can modify the arrangement to obtain a simpler non-sweepable arrangement, thus arriving at a contradiction. If no curve in $\Gamma\setminus\{\alpha,\beta\}$ intersects $L$, then we split the proof into further sub-cases: Since $\Gamma$ was assumed to be a simplest non-sweepable arrangement, bypassing $L$ results in a sweepable arrangement. Thus, there is either a digon/triangle cell on $\gamma$. If the a newly created cell is a digon cell on $\gamma$, we show that removing the curve forming the digon cell results in a simpler non-sweepable arrangement. If we create a triangle cell, we require a slightly more elaborate modification to obtain a simpler non-sweepable arrangement. Thus, in all cases, we obtain a contradiction, and hence there is always either a digon/triangle cell on $\gamma$.
@@
Given any family of 2-intersecting curves $\Gamma$, a curve $\gamma\in\Gamma$ and a point $P\in\tilde{\gamma}$, $\Gamma$ can be swept by $\gamma$ using the following operations: i) passing a digon cell, ii) passing a triangle cell and iii) taking a loop, so that at any point in time during the sweep, the family of curves remain 2-intersecting, and $P$ lies in the interior of $\tilde{\gamma}$.
@@
Suppose the statement is not true. Consider a simplest non-sweepable family $\Gamma$ with sweep curve $\gamma$ i.e., the simplest family for which the statement does not hold. Every curve in this family must intersect $\tilde{\gamma}$ as otherwise we can remove it from $\Gamma$ to obtain a simpler non-sweepable family. If there is a curve $\alpha$ lying entirely in $\tilde{\gamma}$ it cannot be visible from $\gamma$, as otherwise we could take a loop on $\alpha$ (while still keeping $P$ within $\tilde{\gamma}$) to obtain a simpler non-sweepable family - since this decreases the number of regions contained in $\gamma$. If $\alpha$ is not visible from $\gamma$, $\Gamma^{\prime}=\Gamma\setminus\{\alpha\}$ is a simpler non-sweepable arrangement contradicting the minimality of $\Gamma$. If the curves in $\Gamma\setminus\{\gamma\}$ pairwise intersect at most once in $\tilde{\gamma}$, then by Lemma 8, the arrangement is sweepable.
Therefore, in the arrangement $\Gamma$, we can assume that $(i)$ each curve in $\Gamma\setminus\{\gamma\}$ intersects $\gamma$ twice, and there is a pair of curves intersecting twice in $\tilde{\gamma}$, i.e., they form a digon in $\tilde{\gamma}$, $(ii)$ each minimal lens in $\tilde{\gamma}$ is visible from $\gamma$, and $(iii)$ there is either no digon/triangle cell on $\gamma$, or there is exactly one digon/triangle $C$ cell in the arrangement, and $C$ contains the point $P$.
Let $L$ be a minimal digon contained in $\tilde{\gamma}$ and suppose that $L$ is defined by the curves $\alpha,\beta\in\Gamma\setminus\{\gamma\}$ and has vertices $u$ and $v$. $L$ must be visible from $\gamma$, as otherwise by Lemma 4, bypassing $L$ results in a simpler non-sweepable arrangement. Let $a_{1}$ and $a_{2}$ be the intersection points of $\alpha$ with $\gamma$ s.t. the points $a_{1},u,v,a_{2}$ lie in cyclic order along $\alpha$.Similarly, let $b_{1}$ and $b_{2}$ be the intersection points of $\beta$ and $\gamma$ so that $b_{1},u,v,b_{2}$ appear in cyclic order along $\beta$. See Figure 13. Let $u_{\alpha},u_{\beta},v_{\alpha}$ and $v_{\beta}$ be the reference points for bypassing lens $L$. We now split the proof into two cases depending on whether $L$ is non-empty, i.e., there is a curve $\delta\in\Gamma\setminus\{\alpha,\beta,\gamma\}$ that intersects $L$, or is empty, i.e., there is no such curve.
@@
Let $x$ be the other end-point of the segment whose one end-point is $y$. Note that $x$ lies on $\alpha$. Let $d_{1},d_{2}$ be the intersections of $\delta$ and $\gamma$ so that $d_{1},x,y,d_{2}$ appear in cyclic order along $\delta$. The cell $C$ containing $u_{\alpha}$ after bypassing $L$ cannot be a digon cell on $\gamma$ since such a digon cell must be the digon defined by $\alpha$ and $\gamma$ and containing $u_{\alpha}$. However that digon contains the point $y$ and cannot be a cell in the arrangement. $C$ can however be a triangle cell on $\gamma$ and in this case it is defined by the curves $\alpha,\gamma$ and $\delta$ and has vertices $a_{1},y$ and $d_{2}$ (Figure 13(a)). Note that this requires that $d_{2}$ lies on the arc of $\gamma$ between $a_{1}$ and $a_{2}$. Also note that if this is the case then, in the original arrangement (before bypassing $L$), there are no intersection points between $a_{1}$ and $u$ on $\alpha$ and between $u$ and $y$ on $\beta$. This along with the fact that $L$ is minimal implies that triangle $T$ with vertices $x,a_{1}$ and $d_{2}$ is a minimal triangle on $\gamma$. Therefore by Lemma 6 we can bypass it (as shown in Figure 13(b)) so that the resulting arrangement is still $2$-intersecting and has one less vertex (since we lose the intersection $x$) inside $\tilde{\gamma}$. Since the new arrangement is simpler, it must be sweepable which means there must be a digon/triangle cell on $\gamma$ not containing $P$ created as a result of bypassing $T$. By Lemma 7, only the cells containing the reference points $x_{\alpha}$ or $x_{\delta}$ can possibly be a digon/triangle cell on $\gamma$. However, the cell containing $x_{\alpha}$ cannot be on $\gamma$ as it is contained in a lens formed by the modified $\alpha$ and $\beta$ that lies in the interior of ${\gamma}$. Therefore, only the cell containing $x_{\delta}$ could be a digon/triangle cell on $\gamma$. However, this cell can lie on $\gamma$ only if $d_{1}$ lies in the arc of $\gamma$ between $b_{2}$ and $b_{1}$ (see Figure 13(c)). In particular, this implies that the interior of the arc of $\delta$ between $x$ and $d_{1}$ is not intersected by any curve in the original arrangement $\Gamma$. Thus, $\delta$ is the desired curve.
@@
Since $\Gamma$ was not sweepable, all cells on $\gamma$ excluding the cell containing $P$ have at least 4 vertices. Now, consider the arrangement $\Gamma^{\prime}=\Gamma\setminus\{\delta\}$, where $\delta$ is the curve guaranteed by Claim 10. We claim that Γsuperscriptnormal-$\Gamma^{\prime}$ remains non-sweepable. Note that the arc of $\delta$ between $d_{2}$ and $y$ separates two cells $C_{1}$ and $C_{2}$. Either one of them contains $P$, or both cells have at least four vertices. Thus, the new cell created by merging these two cells upon removal of $\delta$ is non-sweepable. An analogous statement holds for the arc of $\delta$ between $d_{1}$ and $x$. Finally note that the remaining new cells created as a result of removing the portion of $\delta$ between $x$ and $y$ are not on cells on $\gamma$ since they lie within the lens $L$. Thus, Γsuperscriptnormal-$\Gamma^{\prime}$ is a simpler non-sweepable arrangement contradicting the minimality of $\Gamma$.
@@
Let $D_{u}$ be the cell in $\Gamma$ contained in $\tilde{\gamma}$ with $u$ as a vertex, and has portions of the arcs $[a_{1},u]$ and $[b_{1},u]$ as two sides. Similarly, let $D_{v}$ be the cell containing vertex $v$, and portions of the arcs $[a_{2},v]$ and $[b_{2},v]$ as sides. By Lemma 5, the only digon/triangle cells on $\gamma$ in the new arrangement can be the cells that contain one of the reference points $u_{\alpha},u_{\beta},v_{\alpha}$, or $v_{\beta}$ (See Figure 7). Since $L$ was empty, there is a single cell containing $u_{\alpha}$ and $v_{\alpha}$. Let $C_{\alpha}$ denote this cell. Similarly, let $C_{\beta}$ denote the cell containing the reference points $u_{\beta}$ and $v_{\beta}$.
@@
We treat $C_{\alpha}$ and $C_{\beta}$ independently. Let $C$ denote one of these regions. If $C$ is non-sweepable, we don’t do anything. If $C$ is a digon cell not containing $P$, we throw away the corresponding curve ($\alpha$ or $\beta$). If $C$ is a triangle cell not containing $P$, we do a more elaborate modification of the arrangement described below.

### v2

@@ 4 Sweeping
    Take a loop: Let $\alpha$ be a curve that does not intersect $\gamma$ and such that $\alpha$ is visible from $\gamma$ via the visibility arc $\tau(u,v)$. We define “taking a loop” as the operation that modifies $\gamma$ to $\gamma^{\prime}$ as follows: Taking two points $u^{\prime},u^{\prime\prime}$ very close to $u$ on either side to $u$ on $\gamma$, we replace the segment of $\gamma$ between $u^{\prime}$ and $u^{\prime\prime}$ by a curve between $u^{\prime}$ and $u^{\prime\prime}$ that lies very close to $\tau(u,v)$ and loops around $v$ crossing $\alpha$ twice. See Figures 2(a) and 2(b). More formally, $\gamma$ is modified to $\gamma^{\prime}$ so that $\tilde{\gamma^{\prime}}=\tilde{\gamma}\setminus(\tau(u,v)\oplus B_{\epsilon})$ where $\oplus$ denotes Minkowski sum and $B_{\epsilon}$ is a ball of radius $\epsilon$ for an arbitrarily small $\epsilon$. In the rest of the paper, we avoid such formal definitions and resort to figures for ease of exposition. However, our informal definitions using figures can easily be formalized.
@@
    Bypass a digon cell: Suppose that a curve $\alpha$ and the sweep curve $\gamma$ form digon cell $D$ with vertices $u$ and $v$. Then, we define the operation of “bypassing $D$” as the modification of $\gamma$ to a curve $\gamma^{\prime}$ that goes around the digon cell $D$ as shown in Figures 3(a) and 3(b).
@@
    Bypass a triangle cell: Let $\alpha$ and $\beta$ be two curves that along with the sweep curve $\gamma$ define a triangle cell $T$. We define the operation of “bypassing $T$” as the modification of $\gamma$ to a curve $\gamma^{\prime}$ which goes around $T$ as show in Figures 4(a) and 4(b).
@@
    Bypass a visible vertex: Let $v$ be a vertex lying inside $\tilde{\gamma}$ where the curves $\alpha$ and $\beta$ intersect. If there is a visibility arc $\tau(u,v)$ from a point $u$ on $\gamma$ to $v$ whose interior lies in the interior of $\tilde{\alpha}$ as well as $\tilde{\beta}$, then by definition, the interior of $\tau(u,v)$ also lies in the interior of $\tilde{\gamma}$. In this case, we refer to $\tau(u,v)$ as the bypassability curve for $v$. “Bypassing $v$” is the modification of $\gamma$ to a curve $\gamma^{\prime}$ that takes a loop around $v$ as shown in Figures 5(a) and 5(b). Note that unlike the previous operations, the orientation of the curves are important here.
@@ 7 A simpler proof of the theorem of Snoeyink and Hershberger
Let $\Gamma$ be an arrangement of 2-intersecting curves with sweep curve $\gamma$, and let $P$ be a specified point in $\tilde{\gamma}$ that does not lie on any of the curves in $\Gamma$. Our goal is to sweep the arrangement of $\Gamma$ with $\gamma$ so that the arrangement remains 2-intersecting throughout the process and $P$ remains inside $\tilde{\gamma}$ until no curve in $\Gamma\setminus\gamma$ intersects $\tilde{\gamma}$. We restrict ourselves to the remaining sweeping operations, namely: $(i)$ taking a loop on a curve that lies entirely in $\tilde{\gamma}$, $(ii)$ bypassing a digon cell, and $(iii)$ bypassing a triangle cell. We do not use the operation of bypassing a visible vertex, as this violates the 2-intersecting property of the curves.
@@
Broad idea. In the restricted setting where each pair of curves in $\Gamma\setminus\{\gamma\}$ intersects at most once inside $\tilde{\gamma}$, we can show that the arrangement is sweepable. Therefore, a simplest non-sweepable arrangement must have at least one pair of curves that intersect twice (i.e., they form a digon) in $\tilde{\gamma}$. Bypassing this digon results in a sweepable arrangement. However, we show that we can modify this resulting arrangement by bypassing a minimal digon, bypassing a minimal triangle (defined in Section 5), and applying a sequence of digon cell bypassing and triangle cell bypassing operations to obtain a simpler non-sweepable arrangement, thus arriving at a contradiction.
@@
Let $\Gamma$ be a set of 2-intersecting curves with sweep curve $\gamma$ s.t. all curves in $\Gamma\setminus\{\gamma\}$ intersect $\gamma$, and any pair of curves in $\Gamma\setminus\{\gamma\}$ intersect at most once in $\tilde{\gamma}$. Let $P$ be any prescribed point in $\tilde{\gamma}$. Then, there is a sweepable cell on $\gamma$, i.e., there is either a digon/triangle cell on $\gamma$ that does not contain $P$.
Proof 7.1.
In order to prove the lemma, we need the following claim which though not explicitly stated is also proved as part of Lemma 3.1 of [23].
@@
Proof 7.2.
We prove by induction on the number of curves intersecting the side of $T$ on $\gamma$. If no curve intersects the interior of this side, since the curves are 1-intersecting in $\tilde{\gamma}$, it follows that $T$ is a triangle cell. Assume that the claim holds for any half-triangle with less than $k$ curves (where $k\geq 1$) intersecting the side of $T$ on $\gamma$. Suppose now that there are $k$ curves intersecting the interior of the side of $T$ on $\gamma$.
Let $x$ denote the intersection point of $\beta$ and $\gamma$ on the boundary of $T$. See Figure 12. Let $v$ denote the intersection point of $\alpha$ and $\beta$ on the boundary of $T$. Since no curve can intersect the side of $T$ on $\alpha$, every curve $\delta^{\prime}$ intersecting the interior of $T$ intersects the interior of the sides $T$ on $\beta$ and $\gamma$ at one point each. Walking from $x$ to $v$ along $\beta$ on the boundary of $T$, let $\delta$ be the first curve intersecting $\beta$. Now, $\beta$ and $\delta$ form a half-triangle $T^{\prime\prime}$ on $\gamma$ with edge $\beta$ and s.t. less than $k$ curves intersect the interior of the side of $T^{\prime\prime}$ on $\gamma$. Since every curve in $\Gamma$ has at most one intersection point on the side of $T$ on $\gamma$, the same holds for $T^{\prime\prime}$. Hence, by the inductive hypothesis, there is a triangle cell $T^{\prime}$ on $\gamma$ that lies in $T^{\prime\prime}$. Since $T^{\prime\prime}$ lies in $T$, $T^{\prime}$ is the claimed triangle cell in $T$.
@@
The proof of the main result in this section (Theorem 7.3) follows the structure of Theorem 1, i.e., proof for non-piercing regions. Unlike in the non-piercing case however, since the curves are 2-intersecting, no pair of curves form a negative lens in $\tilde{\gamma}$. Let $L$ be a minimal lens in a simplest non-sweepable arrangement $\Gamma$ defined by curves $\alpha,\beta\in\Gamma$. We split the proof into two cases. In the first case, there is a curve $\delta\in\Gamma\setminus\{\alpha,\beta\}$ intersecting $L$. In this case, we show that we can modify the arrangement to obtain a simpler non-sweepable arrangement, thus arriving at a contradiction. If no curve in $\Gamma\setminus\{\alpha,\beta\}$ intersects $L$, we split the proof into further sub-cases. Since $\Gamma$ was assumed to be a simplest non-sweepable arrangement, bypassing $L$ results in a sweepable arrangement. Thus, there is either a digon or a triangle cell created on $\gamma$. If the newly created cell is a digon cell on $\gamma$, we show that removing the curve forming the digon cell along with $\gamma$ results in a simpler non-sweepable arrangement. If the newly created cell is a triangle cell, we require a more elaborate modification to obtain a simpler non-sweepable arrangement. Thus, in all cases, we obtain a contradiction.
@@
Given any family of 2-intersecting curves $\Gamma$, a curve $\gamma\in\Gamma$ and a point $P\in\tilde{\gamma}$ not lying on any of the curves in $\Gamma$, $\Gamma$ can be swept by $\gamma$ using the following operations: i) passing a digon cell, ii) passing a triangle cell and iii) taking a loop, so that at any point in time during the sweep, the family of curves remain 2-intersecting, and $P$ lies in the interior of $\tilde{\gamma}$.
@@
Suppose that the statement is not true. Consider a simplest non-sweepable family $\Gamma$ with sweep curve $\gamma$ for which the statement does not hold. Every curve in this family must intersect $\tilde{\gamma}$ as otherwise we can remove it from $\Gamma$ and obtain a simpler non-sweepable family. If there is a curve $\alpha$ lying entirely in $\tilde{\gamma}$ it cannot be visible from $\gamma$, as otherwise we could take a loop on $\alpha$ (while still keeping $P$ within $\tilde{\gamma}$) to obtain a simpler non-sweepable family - since this decreases the number of regions contained in $\gamma$. On the other hand, if $\alpha$ is not visible from $\gamma$, $\Gamma^{\prime}=\Gamma\setminus\{\alpha\}$ is a simpler non-sweepable arrangement contradicting the minimality of $\Gamma$. Thus, there is no curve $\alpha$ lying entirely in $\tilde{\gamma}$.
If the curves in $\Gamma\setminus\{\gamma\}$ pairwise intersect at most once in $\tilde{\gamma}$, then by Lemma 8, the arrangement is sweepable. Therefore, in the arrangement $\Gamma$, we have that $(i)$ each curve in $\Gamma\setminus\{\gamma\}$ intersects $\gamma$ twice, and there is a pair of curves intersecting twice in $\tilde{\gamma}$, i.e., they form a digon in $\tilde{\gamma}$, and $(ii)$ there is either no digon/triangle cell on $\gamma$, or there is exactly one digon/triangle $C$ cell in the arrangement, and $C$ contains the point $P$.
Let $L$ be a minimal digon contained in $\tilde{\gamma}$ and suppose that $L$ is defined by the curves $\alpha,\beta\in\Gamma\setminus\{\gamma\}$ and has vertices $u$ and $v$. $L$ must be visible from $\gamma$, as otherwise by Lemma 4, bypassing $L$ results in a simpler non-sweepable arrangement. Let $a_{1}$ and $a_{2}$ be the intersection points of $\alpha$ with $\gamma$ s.t. the points $a_{1},u,v,a_{2}$ lie in cyclic order along $\alpha$. Similarly, let $b_{1}$ and $b_{2}$ be the intersection points of $\beta$ and $\gamma$ so that $b_{1},u,v,b_{2}$ appear in cyclic order along $\beta$. See Figure 13. Let $u_{\alpha},u_{\beta},v_{\alpha}$ and $v_{\beta}$ be the reference points for bypassing lens $L$. We now split the proof into two cases depending on whether $L$ is non-empty, i.e., there is a curve $\delta\in\Gamma\setminus\{\alpha,\beta,\gamma\}$ that intersects $L$, or it is empty, i.e., there is no such curve.
@@
Let $x$ be the other end-point of the segment whose one end-point is $y$. Note that $x$ lies on $\alpha$. Let $d_{1},d_{2}$ be the intersections of $\delta$ and $\gamma$ so that $d_{1},x,y,d_{2}$ appear in cyclic order along $\delta$. The cell $C$ containing $u_{\alpha}$ after bypassing $L$ cannot be a digon cell on $\gamma$ since such a digon cell must be the digon defined by $\alpha$ and $\gamma$ and containing $u_{\alpha}$. However that digon contains the point $y$ and cannot be a cell in the arrangement. $C$ can however be a triangle cell on $\gamma$ and in this case it is defined by the curves $\alpha,\gamma$ and $\delta$ and has vertices $a_{1},y$ and $d_{2}$ (Figure 14(a)). Note that this requires that $d_{2}$ lies on the arc of $\gamma$ between $a_{1}$ and $a_{2}$. Also note that if this is the case then, in the original arrangement (before bypassing $L$), there are no intersection points between $a_{1}$ and $u$ on $\alpha$ and between $u$ and $y$ on $\beta$. This along with the fact that $L$ is minimal implies that triangle $T$ with vertices $x,a_{1}$ and $d_{2}$ is a minimal triangle on $\gamma$. Therefore by Lemma 6 we can bypass it (as shown in Figures 14(b) and 14(c)) so that the resulting arrangement is still $2$-intersecting and has one less vertex (since we lose the intersection $x$) inside $\tilde{\gamma}$. Since the new arrangement is simpler, it must be sweepable which means there must be a digon/triangle cell on $\gamma$ not containing $P$ created as a result of bypassing $T$. By Lemma 7, only the cells containing the reference points $x_{\alpha}$ or $x_{\delta}$ can possibly be a digon/triangle cell on $\gamma$. However, the cell containing $x_{\alpha}$ cannot be on $\gamma$ as it is contained in a lens formed by the modified $\alpha$ and $\beta$ that lies in the interior of ${\gamma}$. Therefore, only the cell containing $x_{\delta}$ could be a digon/triangle cell on $\gamma$. However, this cell can lie on $\gamma$ only if $d_{1}$ lies in the arc of $\gamma$ between $b_{2}$ and $b_{1}$ (see Figure 14(c)). In particular, this implies that the interior of the arc of $\delta$ between $x$ and $d_{1}$ is not intersected by any curve in the original arrangement $\Gamma$. Thus, $\delta$ is the desired curve.
@@
Since $\Gamma$ was not sweepable, all cells on $\gamma$ excluding the cell containing $P$ have at least 4 vertices. Now, consider the arrangement $\Gamma^{\prime}=\Gamma\setminus\{\delta\}$, where $\delta$ is the curve guaranteed by Claim 10. We claim that $\Gamma^{\prime}$ remains non-sweepable. Note that the arc of $\delta$ between $d_{2}$ and $y$ separates two cells $C_{1}$ and $C_{2}$. Either one of them contains $P$, or both cells have at least four vertices. Thus, the new cell created by merging these two cells upon removal of $\delta$ is non-sweepable. An analogous statement holds for the arc of $\delta$ between $d_{1}$ and $x$. Finally note that the remaining new cells created as a result of removing the portion of $\delta$ between $x$ and $y$ are not cells on $\gamma$ since they lie within the lens $L$. Thus, $\Gamma^{\prime}$ is a simpler non-sweepable arrangement contradicting the minimality of $\Gamma$.
@@
Let $D_{u}$ be the cell in $\Gamma$ contained in $\tilde{\gamma}$ with $u$ as a vertex, and has portions of the arcs $[a_{1},u]$ and $[b_{1},u]$ as two sides. Similarly, let $D_{v}$ be the cell containing vertex $v$, and portions of the arcs $[a_{2},v]$ and $[b_{2},v]$ as sides. By Lemma 5, the only digon/triangle cells on $\gamma$ in the new arrangement can be the cells that contain one of the reference points $u_{\alpha},u_{\beta},v_{\alpha}$, or $v_{\beta}$ (See Figure 7). Since $L$ was empty, after bypassing, there is a single cell containing $u_{\alpha}$ and $v_{\alpha}$. Let $C_{\alpha}$ denote this cell. Similarly, let $C_{\beta}$ denote the cell containing the reference points $u_{\beta}$ and $v_{\beta}$ (after $L$ is bypassed).
@@
We treat $C_{\alpha}$ and $C_{\beta}$ independently. Let $C$ denote one of these regions. If $C$ is non-sweepable, we don’t do anything. If $C$ is a digon cell not containing $P$, we throw away the corresponding curve ($\alpha$ or $\beta$). If $C$ is a triangle cell not containing $P$, we do a more elaborate modification of the arrangement. We start with the case when $C$ is a digon cell.
Case 2a. $C$ is a digon cell. We assume without loss of generality that the cell $C_{\alpha}$ forms a digon cell in the modified arrangement. Then, the cell $D_{u}$ (in the original arrangement $\Gamma$) has the arc $[a_{1},u]$ as one of its sides, and also has one side on $\gamma$. Since $\Gamma$ is non-sweepable, $D_{u}$ either contains $P$ or has at least 4 vertices, of which at least two are not on $\alpha$. Similarly $D_{v}$ either contains $P$ or has at least 4 vertices, of which at least two are not on $\alpha$. Now, consider the arrangement $\Gamma^{\prime}=\Gamma\setminus\{\alpha\}$. Then, the cell in $\Gamma^{\prime}$ that contains the region corresponding to $C_{\alpha}$ either contains $P$ or has at least 4 vertices namely the vertices of $D_{u}$ and $D_{v}$ that are not on $\alpha$. Since $\Gamma^{\prime}$ is simpler than $\Gamma$, it must be sweepable. Since no cell that has a side contributed by $\alpha$ is unchanged, it follows that $C_{\beta}$ is the only sweepable cell in $\Gamma^{\prime}$.
Suppose $C_{\beta}$ is a digon cell, let $\Gamma^{\prime\prime}$ denote the arrangement obtained by removing $\beta$ from $\Gamma^{\prime}$. Then, the cell $C^{\prime\prime}$ in $\Gamma^{\prime\prime}$ that contains the region corresponding to $C_{\beta}$ either has $\geq 4$ sides, or contains $P$. To see this, note that if either $D_{u}$ or $D_{v}$ contain $P$ then $C^{\prime\prime}$ also contains $P$. If neither $D_{u}$ nor $D_{v}$ contain $P$, then both of them must have at least two vertices each that are not on either $\alpha$ or $\beta$ these are also vertices of $C^{\prime\prime}$ implying that $C^{\prime\prime}$ has at least $4$ sides. Thus, $\Gamma^{\prime\prime}$ is a simpler non-sweepable arrangement than $\Gamma$, contradicting our assumptions. See Figure 15. If $C_{\beta}$ is a triangle cell, we handle this in Case 2b..
(a) Either or both of $C_{\alpha}$ or $C_{\beta}$ is a digon cell after bypassing the lens
(b) $C_{\alpha}$ has at least four sides in $\Gamma^{\prime}=\Gamma\setminus\{\alpha\}$.
(c) If one of $D_{u}$ or $D_{v}$ contains the point $P$ then $C_{\alpha}$ also contains $P$ in $\Gamma^{\prime}$
Figure 15: Illustration of case 2a
Case 2b. $C$ is a triangle cell. In this case, we find a minimal triangle formed by one of $\alpha$ or $\beta$ and $\gamma$ along with a third curve $\delta$. But, just bypassing this minimal triangle may create additional bypassable cells on $\gamma$ in $\tilde{\gamma}$. We could bypass these cells and show that we would eventually obtain a simpler non-sweepable arrangement. It is however, easier to deal with such potentially sweepable cells before bypassing a minimal cell involving $\gamma$ so that we immediately obtain a simpler non-sweepable arrangement.